Part A Part complete Calculate [H3O+] in the following aqueous solution at 25 ∘C : [OH−]=...

Part A

Part complete

Calculate [H3O+] in the following aqueous solution at 25 ∘C : [OH−]= 5.3×10−4 M .

Express your answer using two significant figures.

--

SubmitPrevious AnswersRequest Answer

Part B

Part complete

Calculate [H3O+] in the following aqueous solution at 25 ∘C : [OH−]= 2.4×10−12 M .

Express your answer using two significant figures.

--

[H3O+]

[ H 3 O + ]

=

nothing

M

SubmitPrevious AnswersRequest Answer

Part C

Part complete

Calculate [H3O+] in the following aqueous solution at 25 ∘C : [OH−]= 2.2×10−2 M .

Express your answer using two significant figures.

--

[H3O+]

[ H 3 O + ]

=

nothing

M

SubmitPrevious AnswersRequest Answer

Part D

Part complete

Classify the solutions as acidic or basic.

Drag the appropriate items to their respective bins.

--

[H3O+]

[ H 3 O + ]

=

nothing

M

Reset

Help

  • [OH−]= 5.3×10−4 M

    [ O H Superscript-] = 5.3 times 10 Superscript -4 ~ M

  • [OH−]=2.2×10−2 M

    [ O H Superscript-] = 2.2 times 10 Superscript -2 ~ M

  • [OH−]=2.4×10−12 M

    [ O H Superscript-] = 2.4 times 10 Superscript -12 ~ M

  • Acidic solution

  • Basic solution
Answer #1

A) The concentration of H3O + is calculated:

[H3O +] = 10 ^ -14 / [OH-] = 10 ^ -14 / 5.3x10 ^ -4 = 1.89x10 ^ -11 M

B) The concentration of H3O + is calculated:

[H3O +] = 10 ^ -14 / [OH-] = 10 ^ -14 / 2.4x10 ^ -12 = 4.17x10 ^ -3 M

C) The concentration of H3O + is calculated:

[H3O +] = 10 ^ -14 / [OH-] = 10 ^ -14 / 2.2x10 ^ -2 = 4.55x10 ^ -13 M

D) To:

[OH-] = 5.3x10 ^ -4 M

Basic solution

[OH-] = 2.2x10 ^ -2 M

Basic solution

[OH-] = 2.4x10 ^ -12 M

Acid solution

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.